OpenStudy (anonymous):

If the complex number x = 3 + bi and |x|2 = 13, which is a possible value of b? A= 2 B= 4 C= 9 D= 10

OpenStudy (solomonzelman):

take |x|2 = 13 divide both sides by 2, |x|=7.5 7.5 =3 +bi 4.5=bi -4.5i=b or....

OpenStudy (solomonzelman):

|x|=7.5 -7.5 = 3 + bi -10.5=bi +10.5i=b

OpenStudy (solomonzelman):

@ChelsH, do you see how I got this?

OpenStudy (anonymous):

I don't fully understand it..

OpenStudy (anonymous):

The answer doesn't fit one of the answer choices.

OpenStudy (solomonzelman):

I get why, |x|=6.5 not 7.5 my bad

OpenStudy (solomonzelman):